Find the vertices, eccentricity, foci and asymptotes of the hyperbola x2
8
1
Expert's answer
2019-03-11T12:31:03-0400

I don't see equation of the hyperbola in the question. So, I present general way and you can substitute any numbers.

This is standard form of the equation of a hyperbola with center "(h,k)" is

"\dfrac{(x-h)^2}{a^2}-\dfrac{(y-k)^2}{b^2}=1."

The vertices are "(-a,0), \,\, (a,0)".

To find the foci, we need the distance from the center to the foci "c^2=a^2+b^2".

The foci are "(-c,0), (c,0)".

The asymptotes are

"\dfrac{(x-h)^2}{a^2}-\dfrac{(y-k)^2}{b^2}=0 \Rightarrow y=\pm \dfrac{b}{a}(x-h)+k."